Problem 13-6 New-Project Analysis The Campbell Company is considering adding a robotic paint sprayer to its production line. The sprayers base price is $1,080,000, and it would cost another $20,500 to install it. The machine falls into the MACRS 3-year class, and it would be sold after 3 years for $632,000. The MACRS rates for the first three years are 0.3333, 0.4445, 0.1481, and 0.0741. The machine would require an increase in net working capital (inventory) of $20,000. The sprayer would not change revenues, but it is expected to save the firm $399,000 per year in before-tax operating costs, mainly labor. Campbells marginal tax rate is 35%. a. What is the Year-0 net cash flow? If the answer is negative, use minus sign. b. What are the net operating cash flows in Years 1, 2, and 3 Round your answers to the nearest dollar. Year 1 Year 2 Year 3 C. What is the additional Year-3 cash flow (i.e, the after-tax salvage and the return of working capital)? Round your answer to the nearest dollar d. If the projects cost of capital is 15 %, what is the NPV of the project? Round your answer to the nearest dollar. Should the machine be purchased? -Select-
Answer #1
(a) Year 0 net Cash Flow:
A Basic Price $1,080,000
B Installation cost $20,500
E=A+B Depreciable Cost $1,100,500
C Increase in net working capital $20,000
D=A+B+C Year 0 net Cash Flow: $1,120,500
(b) Net Operating Cash flow in Year 1,2,and 3
Year 1 2 3 4
A Before tax savings $             399,000 $        399,000 $        399,000
B Tax Rate                         0.35                    0.35                    0.35
C=A*(1-B) After Tax Saving $             259,350 $        259,350 $        259,350
D MACRS Recovery Rate 0.3333 0.4445 0.1481 0.0741
E=1100500*D Amount of depreciation $             366,797 $        489,172 $        162,984 $          81,547
F=E*B Depreciation Tax Shield $             128,379 $        171,210 $           57,044
G=C+F Net Operating cash flow $             387,729 $        430,560 $        316,394
Net Operating Cash flow in Year 1,2,and 3
Year1 $             387,729
Year2 $             430,560
Year3 $             316,394
.(c) Additional year 3 Cash flow
A Book value of Equipment at end of year3 $                81,547
B Selling Price after 3 years $632,000
C=B-A Gain on disposal $550,453
D=C*0.35 Taxes on Gain $192,658.53
E=B-D Cash flow due to disposal $439,341.47
F Release of Working Capital $20,000
G=E+F Additional year 3 Cash flow $459,341.47
(d) Calculation of Net Present Value(NPV)
Present Value(PV) of Cash flow=(Cash Flow)/((1+i)^N)
i=discount rate=cost of capital=15%=0.15
N=Year of Cash flow
Net Present Value (NPV)=Sum of PV of Cash flows
N Year 0 1 2 3
A Initial cash flow ($1,120,500)
B Net Operating cash flow $        387,729 $        430,560 $        316,394
C Additional Cash flow end of year 3 $459,341.47
D Net Cash Flow ($1,120,500) $        387,729 $        430,560 $        775,736 SUM
E=D/(1.15^N) Present Value of Cash Flows $ (1,120,500.00) $ 337,155.50 $ 325,565.43 $ 510,058.94 $ 52,279.87
Sum of Present Values of Cash flows $          52,279.87
NPV of the project $          52,279.87
YES , The Machine should be Purchased
NPV positive
Hence, Rate of return is higher than cost of capital of 15%